# Multi: Multimodal Understanding Leaderboard with Text and Images

### v2

@@ 3.4 Comparison with Existing Benchmarks
MMBench Liu et al. (2023c) en 20 3K - 3K - Web, Repurposed
SEED-Bench Li et al. (2023b) en 12 19K - 19K+ - Anno.
SEED-Bench-2 Li et al. (2023a) en 27 24K - 22K+ - Anno.
@@
Table 1: The comparison between Multi and other existing benchmarks. Sub: Subject or Field, Q: Question, Ana: Analysis or Explanations, Img: Images, Kn: Knowledge or Lecture. NI: the question with pure text, SI: the question with a single image, MI: the question with multiple images. SA: multiple-choice question with single correct answer, MA: multiple-choice question with multiple correct answers, FB: fill-in-the-blank question (no more than 10 words), OP: open-ended writing question (more than 10 words). Anno.: Annotation
Multi demonstrates a comprehensive blend of features that surpasses existing benchmarks in several dimensions. Notably, Multi covers a wide array of subjects and a substantial number of questions (18K), as well as over 10K analysis and 4.6K extensive knowledge content, which is considerably larger than most benchmarks, ensuring a broad and diverse testing environment. Multi possesses 7.7K images, which is essential for benchmarking MLLMs that require visual understanding alongside textual information. The inclusion of both single and multiple image questions, as well as a variety of answer types, makes Multi a versatile and challenging benchmark. Furthermore, the questions without images also test the MLLMs’ ability on dealing with plain text information. Meanwhile, the various sources, complex annotation, and processing stages provide sufficient augmentation to alleviate data leakage. Multi not only encompasses variations of classic questions but also includes recently updated questions, which significantly enhances its diversity.
@@ 4 Experiments
4.1 Models
We evaluate a wide range of MLLMs that support Chinese, including Chinese-LLaVA LinkSoul-AI (2023), Qwen-VL Bai et al. (2023a), VisCPM Hu et al. (2023), VisualGLM Du et al. (2022), InternVL Chen et al. (2023), Yi-VL 01.ai (2023), Gemini Vision Team (2023), and GPT-4V OpenAI (2023b). We evaluate these models with both multimodal input and text-only input to verify the information gain of input images. We also select several most capable LLMs for comparison with text-only input, including DFM-2 Chen et al. (2022), MOSS Sun et al. (2023b), ChatGPT OpenAI (2022), GPT-4 OpenAI (2023a), and Gemini Team (2023), and the performance of these models on questions with images will reflect their abilities on finding loss of information. Model specifications are listed in Table 10. Due to the API request rate limit of Gemini and GPTs, ablation studies are mostly performed on weight-accessible models. We choose the checkpoints with the largest model size and the latest version, and use FP16 or INT4 quantization to accelerate inference if officially provided. We follow the official guidelines to prompt each model so that the outputs go in the desired way.
@@ 4.2 Settings
We use specialized prompts for each question, an example shown in Figure 4. The prompts are designed carefully according to the features of each type of question and the answer patterns expected. We also modify the input format to fit into official inference guidelines. The complete collection of prompts is presented in Appendix C.
Figure 4: An example of the prompts used when evaluating a multiple-choice question with image context, knowledge piece and single correct answer.
@@
For pure-text questions, we use the text as input. For some models like VisCPM, InternVL and Yi-VL which compulsorily demand an image in each turn, we feed the model a blank image with color set to RGB(0,0,0) along with plain text in evaluation. For efficiency, results of GPT-4 and Gemini on pure-text questions are directly used as the results of GPT-4V and Gemini Vision respectively.
@@ 4.3 Metrics
We define the total points of an MA question as the number of correct choices, and each correct choice selected is rewarded one point. If the given answer contains any wrong choice, the score will be counted to zero. We report the score ratio (# points / # total points) as the metric. We also report accuracy as a more rigorous metric, where only correctly giving all the choices without wrong ones will be granted points.  [^4] For example, a question with correct answer ACE worth 3 points, and answer AC will be granted 2 points and answer BC or ABCE will be granted 0 points. However, on calculating accuracy none will be counted, and only ACE will be calculated as correct.
@@
We define the total points of a blank-filling question as the number of the blanks marked as [MASK]. It is required in prompts that each line of the given answer corresponds to a blank in order. We follow the most strict standard of exact match. Therefore, only answers exactly matching the standard answers will be granted points. We report the score ratio as the final metric.
@@
The points and counting method is similar to FB, but we use a loose standard and report normalized ROUGE-L Lin (2004) score for each point. Please note that the reference answer may be concise or in detail, and there could be other possible answers.
@@ 4.4 Main Experiment Results
We report the overall and field-specific performance of tested models on the whole benchmark in Table 2, 3, and 4.
Model Overall NI SI MI
Puretext (LLM)
MOSS 32.6 36.1 27.3 17.1
DFM-2.0 49.7 63.0 28.7 11.3
Gemini 52.2 62.5 36.2 18.3
ChatGPT 35.9 54.0 6.8 5.1
GPT-4 50.2 74.5 11.3 8.8
Text+Image (MLLM)
Chinese-LLaVA 28.5 32.3 22.6 17.8
VisualGLM 31.1 35.1 25.2 9.7
VisCPM 33.4 36.8 28.4 16.6
Qwen-VL 39.0 43.2 32.7 20.7
InternVL 44.9 50.9 35.5 25.1
Yi-VL 55.3 63.8 42.0 24.5
Gemini Vision 53.7 62.5 40.0 24.5
GPT-4V 63.7 74.5 46.9 28.1
Table 2: The main performance of models evaluated on Multi. NI: the question with no image, SI: the question with a single image, MI: the question with multiple images.
Overall comparison.
We report the overall performance in Table 2. The most powerful competitor, GPT-4V, achieves a mere 63.7% score, underscoring the benchmark’s complexity and challenge. Yi-VL outperforms other open-source models, but there still remains a notable gap with GPT-4V, and those smaller models do not get as much as half of the scores.
Comparison by number of images.
@@
In Table 2, we also present the performance categorized by image number. For MLLMs, a higher score on the Non-Image (NI) set suggests improved performance on multimodal questions, including the Single Image (SI) set and Multiple Image (MI) set. It is evident that questions requiring more images are more challenging. A significant drop in performance is observed when answering questions with more than one image. Only GPT-4V (28.1%) manages to exceed the average baseline set by random guessing.
Conversely, for LLMs, there exists a reverse correlation between scores on the NI set and those on the SI and MI sets. This is because we prompt the model to determine whether visual information is necessary for answering a question and if so the model needs to refuse to answer. Less capable models may simply make a guess, but more sophisticated models tend to withhold an answer, resulting in lower but more reliable overall scores. The results on SI and MI sets for LLMs indicate a long way before mitigating hallucination.
@@
Puretext (LLM)
MOSS 38.5 33.1 6.8 2.7 8.7
DFM-2.0 55.8 53.9 29.7 13.3 10.3
Gemini 58.2 52.7 22.8 29.1 7.9
ChatGPT 40.0 39.4 17.9 10.5 7.7
GPT-4 51.3 60.0 53.1 32.9 6.8
Text+Image (MLLM)
Chinese-LLaVA 34.5 26.9 3.9 2.4 8.4
@@
InternVL 56.4 33.4 2.1 14.2 13.1
Yi-VL 61.3 42.0 36.4 14.6 8.9
Gemini Vision 59.4 54.4 24.3 30.5 12.5
@@
Table 3: Performance of models on each type of questions of Multi. MA Acc.: Accuracy of MA questions.
Comparison by question type.
In Table 3, we present the performance categorized by question type. A majority of the models achieve their highest scores on the Single Answer Choosing (SA) set, with lower performance on the Multiple Answers Choosing (MA) set. A notable discrepancy is observed between the scores for the MA set and its accuracy, highlighting the smaller models’ inability to identify all correct options accurately.
For the Fill-in-the-Blank (FB) set, which requires short but exact matches, the scores further decline. This is partially due to failure to follow the specified instructions, often leading to correct responses being presented in an unacceptable format.
Furthermore, we note significantly lower scores on the Open-ended Writing (OP) set in comparison to the FB set. VisCPM stands out but only with the best score of 14.1% on the OP set, suggesting that our dataset minimizes the risk of data leakage and poses considerable challenges for models in generation across modalities.
Model JuH SeH Uni Driv AAT
Puretext (LLM)
MOSS 21.2 26.7 23.8 44.1 25.5
DFM-2.0 42.3 42.5 35.7 66.3 3.9
Gemini 47.7 42.3 41.4 66.9 22.5
ChatGPT 31.6 23.7 34.9 52.1 1.3
GPT-4 49.2 33.7 55.1 69.9 0.9
Text+Image (MLLM)
Chinese-LLaVA 21.1 25.4 20.7 35.8 21.8
VisualGLM 22.2 25.6 23.6 40.9 24.9
VisCPM 25.2 28.1 23.0 43.4 23.7
Qwen-VL 32.6 32.9 27.2 49.3 26.4
InternVL 39.3 36.5 30.6 57.7 24.8
Yi-VL 46.6 46.0 45.4 71.1 26.5
Gemini Vision 48.2 45.2 41.7 67.4 27.0
GPT-4V 58.5 52.9 59.0 80.1 26.2
Table 4: Performance of models on each subject of Multi. JuH: level of Junior High school, SeH: level of Senior High school, Uni: level of University, Driv: Chinese driving test, AAT: Administrative Aptitude Test.
Comparison by education level and subjects.
In Table 4, we present the performance categorized by educational levels and subjects. The performance trends for high school and university level questions remain consistent with the overall results observed. For questions at the society level, we anticipate higher scores on the Driving Test. This may be caused by a larger percentage of judgmental questions (in the format of SA with two options), as well as its nature with knowledge of regulations.
Furthermore, questions from the Administrative Aptitude Test (AAT), which typically include at least one image and often examine skills on image pattern recognition (illustrated in the first two examples in the left column of Figure 9), tend to have scores around or below randomly choosing baseline. Even the strongest competitor, GPT-4V, shows limited success, with a performance of only 27.0% on these questions as detailed in the study cited in the paper OpenAI (2023b). This underscores the significant challenge posed by multimodal questions. Notably, the stronger LLMs, specifically DFM-2.0 and the text-only versions of GPT perform poorly on AAT questions as expected, as they often reject answering the majority of them.
4.5 Ablation Study on Image Information Gain
Model NI SI MI
w/o. image w. caption w. ocr w. image w/o. image w. caption w. ocr w. image
Puretext (LLM)
MOSS 36.1 27.3 27.3 (+0.0) 27.6 (+0.3) - 17.1 20.7 (+3.6) 19.0 (+1.9) -
ChatGPT 54.0 6.8 9.9 (+3.1) 6.6 (-0.2) - 5.1 10.7 (+5.6) 5.5 (+0.4) -
Text+Image (MLLM)
Chinese-LLaVA 32.3 26.1 26.3 (+0.2) 25.5 (-0.6) 22.6 (-3.5) 17.6 19.9 (+2.3) 19.6 (+2.0) 17.8 (+0.2)
VisualGLM 35.1 20.8 21.4 (+0.6) 20.4 (-0.4) 25.2 (+4.4) 15.3 15.1 (-0.2) 14.5 (-0.8) 9.7 (-5.6)
VisCPM 36.8 27.1 27.6 (+0.5) 27.2 (+0.1) 28.4 (+1.3) 24.8 21.6 (-3.2) 20.9 (-3.9) 16.6 (-8.2)
Qwen-VL 43.2 30.7 30.3 (-0.4) 31.0 (+0.3) 32.7 (+2.0) 25.5 25.0 (-0.5) 26.2 (+0.7) 20.7 (-4.8)
InternVL 50.9 33.4 33.3 (-0.1) 33.1 (-0.3) 35.5 (+2.1) 24.8 21.9 (-2.9) 22.9 (-1.9) 25.1 (+0.3)
Gemini/Vision 62.5 36.2 36.9 (+0.7) 38.4 (+2.2) 40.0 (+3.8) 18.3 23.2 (+4.9) 18.6 (+0.3) 24.5 (+6.2)
Yi-VL 63.8 39.9 38.7 (-1.2) 39.4 (-0.5) 42.0 (+2.1) 24.1 26.5 (+2.4) 24.2(+0.1) 24.5 (+0.4)
GPT-4/V 74.5 11.3 9.7 (-1.6) 1.9 (-9.4) 46.9 (+35.6) 8.8 9.4 (+0.6) 3.1 (-5.7) 28.1 (+19.3)
average +0.30 -0.35 +5.98 +1.54 -0.30 +0.98
Table 5: Performance of models evaluated on the image set of Multi.
To assess the necessity of images in Multi for solving problems, we conduct an ablation study where we either remove images from the SI and MI sets or substitute them with textual descriptions, such as captions and OCR-derived text. We utilize BLIP2 Li et al. (2023c) for generating image captions and EasyOCR [^5] https://pypi.org/project/easyocr/ to extract text from images. The results are shown in Table 5.
For questions that incorporate a single image (as indicated in the SI column), the presence of images significantly aids in answering the questions, with an average performance boost of 5.98%. Notably, GPT-4V experiences a substantial increase of 35.6% in performance, primarily due to its tendency to abstain from answering in the absence of images.
In settings where images are omitted and replaced by their textual descriptions (captions or OCR text), there’s a marginal improvement of 0.30% observed with captions, but a minor reduction of -0.35% with OCR text. Captions, which generally summarize the images, introduce bilingual elements to the models and usually miss details. OCR text, while detailed, lacks spatial information and is not universally applicable, as some images contain no text at all. Both forms of textual information lower the models’ refusal rate, and LLMs benefit more from these than MLLMs. However, they potentially complicate reasoning processes. Nevertheless, a generic caption is found to be more beneficial than scattered OCR fragments.
For questions that involve multiple images (as discussed in the MI column), we categorize models into three groups: 1) Close-source models, specifically GPT-4V and Gemini Pro, which leverage all images and achieve significant improvement. 2) Open-source models capable of handling multiple images within a dialogue or at a time, namely VisualGLM, VisCPM, and Qwen-VL, all of which exhibit a notable performance decline. 3) Open-source models without multi-image support, like Chinese-LLaVA, InternVL, and Yi-VL, show slight improvements. The second group’s decline could be attributed to their inability to utilize conversation history effectively and remember previously seen images. The third group’s limitation likely stems from providing only the first image, insufficient for comprehending all necessary information to answer the question, but to some degree avoiding distraction.
4.6 Evaluation on Multi-Elite

### v3

@@ 3.5 Comparison with Existing Benchmarks
CMMU He et al. (2024) zh 7 4K - 4K - Exams
GAOKAO-MM Zong and Qiu (2024) zh 8 0.6K - 0.9K - Exams
MMBench Liu et al. (2025) zh/en 20 3K - 3K - Web, Repurposed
SEED-Bench Li et al. (2024a) en 12 19K - 19K+ - Anno.
SEED-Bench-2 Li et al. (2024b) en 27 24K - 22K+ - Anno.
@@
Table 3: The comparison between Multi and other existing knowledge-based benchmarks. Sub: Subject or Field, Q: Question, Ana: Analysis or Explanations, Img: Images, Kn: Knowledge or Lecture. NI: the question with pure text, SI: the question with a single image, MI: the question with multiple images. SA: multiple-choice question with single correct answer, MA: multiple-choice question with multiple correct answers, FB: fill-in-the-blank question (no more than 10 words), OP: open-ended writing question (more than 10 words). Anno.: Annotation.
Multi demonstrates a comprehensive blend of features that surpasses existing benchmarks in several dimensions. Notably, Multi covers a wide array of subjects and a substantial number of questions (18K), as well as over 10K analysis and 4.6K extensive knowledge content, which is considerably larger than most benchmarks, ensuring a broad and diverse testing environment. Multi possesses 7.7K images, which is essential for benchmarking MLLMs that require visual understanding alongside textual information. The inclusion of both single and multiple image questions, as well as a variety of answer types, makes Multi a versatile and challenging benchmark. Furthermore, the questions without images also test the MLLMs’ ability to deal with plain text information. Meanwhile, the various sources, complex annotation, and processing stages provide sufficient augmentation to alleviate data leakage. Multi not only encompasses variations of classic questions but also includes recently updated questions, which significantly enhances its diversity.
@@ 4 Experiments
In this section, we conduct a comprehensive experiment and several ablation studies. We firstly introduce baselines, settings, and metrics in § 4.1, § 4.2 and § 4.3. The main experiment analysis is in § 4.4, followed by several research questions and ablation study results from § 4.5 to § 4.10. We conclude our findings and thoughts as take-home messages in § 4.11
4.1 Baselines
Large Models.
We evaluate a wide range of leading-edge MLLMs that support Chinese, including both open-sourced models LinkSoul-AI (2023); Bai et al. (2023a); Wang et al. (2024); Hu et al. (2023); Du et al. (2022); Chen et al. (2024); 01.ai (2023) and closed-source models OpenAI (2024a, 2023b); Team (2023); Anthropic ; OpenAI (2024b). We evaluate these models with both multimodal input and text-only input to verify the information gained from input images. We also select several most capable and representative LLMs Chen et al. (2022); Sun et al. (2023); OpenAI (2022, 2023a); qwe (2024); Zhang et al. (2023); Cai et al. (2024) for comparison. We also add ablation studies to verify the improvement by giving the OCR information or captions to these LLMs. For detailed information about all models evaluated in Multi, please refer to Table 10 in B.
Human Experts.
Seven human experts from top universities in China are involved in the human evaluation. The complete human evaluation set includes all 500 questions from Multi-Elite and another 500 randomly sampled questions from Multi. Each question was assigned to two experts based on the subjects they had chosen in Gaokao. We follow the rule in MMMUYue et al. (2024) where the experts are allowed to consult their textbooks and Multi-Extend but are forbidden to search the Internet. We calculate the score of Multi as the weighted average of each subject.
Random Baseline.
We also provide a random baseline and a most frequent baseline. For the random baseline, we randomly choose one choice in SA, randomly select/drop each choice in MA, and randomly choose from [A, B, C, D, 1, 2, 3, 4, True, False] in OP, and simply take the question content as the answer.
Figure 5: An example of the prompts used when evaluating a multiple-choice question with image context, knowledge piece, and single correct answer.
@@ 4.2 Settings
We use specialized prompts for each question, an example shown in Figure 5. The prompts are designed carefully according to the features of each type of question and the answer patterns expected. We also modify the input format to fit into official inference guidelines. The complete collection of prompts is presented in C.
@@
For pure-text questions, we use the text as input. For some models like VisCPM, InternVL-1.1, and Yi-VL-34B which compulsorily demand an image in each turn, we feed the model a blank image with color set to RGB(0,0,0) along with plain text in evaluation.
@@ 4.3 Metrics
We follow the metric used in Gaokao and define the total points of an MA question as the number of correct choices, and each correct choice selected is rewarded one point. If the given answer contains any wrong choice, the score will be counted to zero. We report the score ratio (# points / # total points) as the metric. We also report accuracy as a more rigorous metric, where only correctly giving all the choices without wrong ones will be granted points. For example, a question with a correct answer ACE worth 3 points, and the incomplete answer AC will be granted 2 points and the answer BC or ABCE will be granted 0 points. However, on calculating accuracy none will be counted, except for ACE which will be calculated as correct.
@@
We define the total points of a blank-filling question as the number of the blanks marked as [MASK]. It is required in prompts that each line of the given answer corresponds to a blank in order. We follow the most strict standard of exact match. Therefore, only answers exactly matching one of the standard answers will be granted points.
@@
The points and counting method is similar to FB, but we use a loose standard and report normalized ROUGE-L Lin (2004) score for each point. Please note that since the reference answer may be concise or in detail, and there could be other possible answers, the lower score ratio of OP questions is expected and reasonable.
We report the score ratio of all questions as the final metric.
@@ 4.4 Main Experiment Results
In this sub-section we report the performance on Multi as detailed in Table 4, 5, and 6 along with analysis below:
Model Overall JuH SeH Uni Driv AAT
random 24.8 14.0 17.2 20.4 35.8 24.6
most 28.4 14.2 19.7 20.5 42.8 23.3
expert 86.1 87.8 82.9 88.2 91.7 76.7
VisualGLM 31.1 22.2 25.6 23.6 40.9 24.9
VisCPM 33.4 25.2 28.1 23.0 43.4 23.7
Chinese-LLaVA 28.5 21.1 25.4 20.7 35.8 21.8
Qwen-VL 39.0 32.6 32.9 27.2 49.3 26.4
Yi-VL-34B 55.3 46.6 46.0 45.4 71.1 26.5
InternVL-1.1 44.9 39.3 36.5 30.6 57.7 24.8
InternVL2-8B 67.9 63.7 58.9 59.6 83.8 26.0
MiniCPM-V-2.6 55.1 46.3 42.8 38.0 73.7 27.4
Qwen2-VL-7B 68.6 60.4 57.8 50.8 87.5 28.7
Qwen2-VL-72B 76.9 66.3 69.5 69.9 94.3 34.0
GPT-4V 63.7 58.5 52.9 59.0 80.1 26.2
Gemini Vision Pro 53.7 48.2 45.2 41.7 67.4 27.0
GPT-4o-mini 58.2 51.7 43.9 49.6 77.1 25.8
GPT-4o (0806) 69.9 65.6 58.6 62.2 87.1 28.0
Gemini 1.5 Flash 64.9 63.2 56.0 52.6 78.7 26.7
Claude 3.5 Sonnet 57.2 38.7 42.3 26.1 83.3 28.9
GPT-4o (1120) 70.0 65.9 59.0 60.2 86.9 28.2
Gemini 1.5 Pro 71.8 71.6 64.3 65.3 84.2 29.3
Table 4: The main performance of all models evaluated on Multi. JuH: Level of Junior High School, SeH: Level of Senior High School, Uni: Level of University, Driv: Chinese Driving Test, AAT: Administrative Aptitude Test. We categorized the models into two groups based on whether they are accessed through model weights or via an API service. Models in each group are sorted by their release date. The highest scores in each group are bolded, while the second highest is underlined. The scores higher than the human expert baseline are highlighted.
Most models are struggled with Multi.
By the beginning of 2024, GPT-4V achieves a score of 63.7%, while by the end of 2024, the most powerful competitor, Qwen2-VL-72B, achieves a score of 76.9%, underscoring Multi’s complexity and challenge even after near a year of its first release. But more importantly, we observe a huge improvement in performance for open-sourced models like QwenVL and InternVL series, indicating the gap between open-source MLLMs and closed ones is becoming increasingly narrow. However, the gap between these models and human experts remains significant.
Analysis from the perspective of education level and subjects:
In Table 4, we also present the performance categorized by educational levels and subjects. The performance trends for non-society level questions remain consistent with the overall results observed, while the performance of these three levels is similar.
For questions at the society level, i.e. Driv. and AAT, we anticipate higher scores on the Driving Test. This may be caused by a larger percentage of judgmental questions (in the format of SA with two options), as well as its nature with knowledge of regulations. We are happy to say that Qwen2-VL-72B has surpassed the human expert baseline in this subject.
Furthermore, questions from the Administrative Aptitude Test (AAT), which typically include at least one image and often examine skills in image pattern recognition (illustrated in the first two examples in the left column of Figure 14), tend to have scores around or below randomly choosing baseline. Even the strongest competitor, Qwen2-VL-72B, shows limited success, with a performance of only 34.0% on these questions. This underscores the significant challenge posed by multimodal questions.
Analysis from the perspective of different numbers of image:
@@
In Table 5, we present the performance categorized by image number. While the highest score on the None-image set achieves 86.9% by Qwen2-VL-72B, a significant drop in performance is observed when answering questions with more than one image. Only Gemini 1.5 Pro manages to get a 50.0% score while other strong models are around 40-41%. It is evident that questions requiring more images are more challenging. The high performance on the NI set may also caused by the large portion of easy-driving questions as mentioned before.
Analysis from the perspective of different question types.
In Table 6, we present the performance categorized by question type. A majority of the models achieve their highest scores on the Single Answer Choosing (SA) set, with lower performance on the Multiple Answers Choosing (MA) set. A notable discrepancy is observed between the scores for the MA set and its accuracy, highlighting the smaller models’ inability to identify all correct options accurately. While Qwen2-VL-72B once again beat human expert baseline.
For the Fill-in-the-Blank (FB) set, which requires short but accurate matches, the scores further decline. This is partially due to failure to follow the specified instructions, often leading to correct responses being presented in an unacceptable format. Gemini 1.5 and GPT-4o models are more capable in FB questions and surpass Qwen2-VL-72B by large.
Furthermore, we note significantly lower scores on the Open-ended Writing (OP) set in comparison to the FB set. None of the models gets even half of the 46.4 % human expert baseline on the OP set. This also suggests that Multi minimizes the risk of data leakage during its construction and poses considerable challenges for models in generation across modalities.
Model NI SI MI
random 27.2 21.0 16.9
most 32.5 22.0 17.7
expert 86.6 85.3 81.8
VisualGLM 35.1 25.2 9.7
VisCPM 36.8 28.4 16.6
Chinese-LLaVA 32.3 22.6 17.8
Qwen-VL 43.2 32.7 20.7
Yi-VL-34B 63.8 42.0 24.5
InternVL-1.1 50.9 35.5 25.1
InternVL2-8B 78.7 51.6 28.7
MiniCPM-V-2.6 63.8 41.8 24.0
Qwen2-VL-7B 78.1 53.7 34.2
Qwen2-VL-72B 86.9 61.5 41.1
GPT-4V 74.5 46.9 28.1
Gemini Vision Pro 62.5 40.0 24.5
GPT-4o-mini 67.7 43.0 30.4
GPT-4o (0806) 80.0 54.1 40.9
Gemini 1.5 Flash 73.8 50.9 40.0
Claude 3.5 Sonnet 68.3 39.7 23.1
GPT-4o (1120) 79.6 54.9 40.9
Gemini 1.5 Pro 80.2 58.5 50.0
Table 5: Performance on each image type of Multi.
@@
random 30.8 22.5 6.5 0.9 3.4
most 30.7 36.3 21.0 1.6 3.4
expert 91.4 81.1 70.3 83.4 46.4
@@
Yi-VL-34B 61.3 42.0 36.4 14.6 8.9
InternVL-1.1 56.4 33.4 2.1 14.2 13.1
InternVL2-8B 72.6 74.2 64.0 40.9 17.5
MiniCPM-V-2.6 64.7 53.2 31.3 15.2 12.3
Qwen2-VL-7B 75.0 72.7 62.1 35.0 18.5
Qwen2-VL-72B 83.5 84.7 78.1 35.1 19.5
@@
Gemini Vision Pro 59.4 54.4 24.3 30.5 12.5
GPT-4o-mini 65.8 56.6 46.3 30.3 12.4
GPT-4o (0806) 74.5 74.6 63.1 47.6 17.2
Gemini 1.5 Flash 68.4 68.5 52.1 50.9 14.1
Claude 3.5 Sonnet 63.0 64.9 52.9 16.6 9.4
GPT-4o (1120) 74.4 74.2 63.0 49.4 17.7
Gemini 1.5 Pro 74.3 76.8 64.9 62.6 14.6
Table 6: Performance on each type of questions of Multi. MA Acc.: Accuracy of MA questions.
4.5 Is Multi-Elite more challenging?
We conduct evaluations on Multi-Elite, as outlined in Table 7, which includes 500 specifically chosen questions. These questions are selected based on pre-annotated quality and difficulty scores, in addition to the evaluation results on Multi discussed in § 4.4. The selection aims to ensure a more balanced distribution but also picks those high-quality questions that bring challenge to strong MLLMs.
Qwen2-VL-72B and Gemini 1.5 Pro achieve the highest score on Multi-Elite with 53.1% and 46.9%, while scores for other models vary between 10.5% and 40.0%. This highlights the substantial challenge presented by Multi-Elite, indicating significant potential for improvement in tackling extremely difficult questions that require in-depth image understanding and intricate reasoning across modalities.
It is important to highlight the accuracy of multiple answers choosing (MA Acc.) as the most demanding task for MLLMs, necessitating a thorough grasp of the relationships between the choices and the questions, and reflecting model reliability of selecting all answers correctly. Meanwhile, the FB set has the largest gap between the highest model performance and human expert baseline, indicating a great challenge for models to answer accurately in these questions.